# Denoising Diffusion Variational Inference: Diffusion Models as Expressive Variational Posteriors

### v4

@@ 1 Introduction
in which we approximate the marginal likelihood $\log\;p_{\boldsymbol{\theta}}(\mathbf{x})$ of a latent variable model $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})$ with an evidence lower bound $\text{ELBO}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ that is a function of an approximate variational posterior $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$. We assume that $p_{\boldsymbol{\theta}}$ factorizes as $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ and admits efficient sampling: examples of such $p_{\boldsymbol{\theta}}$ include Bayesian networks, topic models (Blei, Ng, and Jordan 2003), variational autoencoders (VAEs), and broad classes of $p_{\boldsymbol{\theta}}$ defined via modern probabilistic programming frameworks (Gordon et al. 2014).
@@
Maximizing $\text{ELBO}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ over ${\boldsymbol{\phi}}$ yields a variational posterior $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ that approximates $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$ as well as a tight bound on $\log\;p_{\boldsymbol{\theta}}(\mathbf{x})$ that serves as a learning objective for $p_{\boldsymbol{\theta}}$. The approximation gap $\log\;p_{\boldsymbol{\theta}}(\mathbf{x})-\max_{\boldsymbol{\phi}}\text{ELBO}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ equals precisely $\min_{\boldsymbol{\phi}}\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}))$, which motivates the design of expressive classes of posteriors $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ that reduce this gap. Recent efforts leverage modern generative models—including normalizing flows (Rezende and Mohamed 2015; Kingma et al. 2016) and generative adversarial networks (Goodfellow et al. 2014; Makhzani et al. 2015)—as expressive model families for $q_{\boldsymbol{\phi}}$ that tighten the ELBO.
@@
This work seeks to further improve variational inference via expressive posteriors based on diffusion models (Ho, Jain, and Abbeel 2020; Song et al. 2020). Diffusion methods have become the de-facto standard for high-quality image synthesis (Rombach et al. 2022; Gokaslan et al. 2024). Here, we use diffusion in latent space to parameterize $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$. We train this distribution with a denoising diffusion-like objective that does not involve adversarial training (Makhzani et al. 2015) or constrained invertible normalizing flow architectures (Kingma et al. 2016). Samples from $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ are obtained via iterative refinement of $\mathbf{z}$, starting from a Gaussian distribution, and gradually forming one that is multi-modal and complex.
[…]
@@
Specifically, we propose denoising diffusion variational inference (DDVI), an approximate inference algorithm defined by a class of approximate posterior distribution based on diffusion and a learning objective inspired by the wake-sleep algorithm (Hinton et al. 1995) that implements regularized variational inference. We also derive extensions of our method to semi-supervised learning and clustering.
@@
We evaluate DDVI on synthetic benchmarks and on a real problem in biological data analysis—inferring human ancestry from genetic data. Our method outperforms strong baselines on 1000 Genomes dataset (Siva 2008) and learns a low-dimensional latent space that preserves biologically meaningful structure (Haghverdi, Buettner, and Theis 2015).
@@ 2 Background
which serves as a tractable surrogate for the marginal log-likelihood (MLL). The gap between the MLL and the ELBO equals precisely $\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}))$—thus, a more expressive $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ may better fit the true posterior and induce a tighter ELBO (Kingma and Welling 2013).
Expressive variational posteriors can be formed by choosing more expressive model families—including auxiliary variable methods (Maaløe et al. 2016), MCMC-based methods (Salimans, Kingma, and Welling 2015), normalizing flows (Rezende and Mohamed 2015)—or improved learning objectives—e.g., adversarial or sample-based losses (Makhzani et al. 2015; Zhao, Song, and Ermon 2017; Si, Bishop, and Kuleshov 2022; Si et al. 2023).
The wake-sleep algorithm (Hinton et al. 1995) optimizes an alternative objective
@@
A diffusion model is defined via a user-specified noising process $q$ that maps data $\mathbf{x}_{0}$ into a sequence of $T$ variables $\mathbf{y}_{1:T}=\mathbf{y}_{1},...,\mathbf{y}_{T}$ that represent increasing levels of corruption to $\mathbf{x}_{0}$. We obtain $\mathbf{y}_{1:T}$ by applying a Markov chain $q(\mathbf{y}_{1:T}|\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{y}_{t}|\mathbf{y}_{t-1})$, where we define $\mathbf{y}_{0}=\mathbf{x}_{0}$ for convenience. When $\mathbf{x}_{0}$ is a continuous vector, a standard choice of transition kernel is $q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\sqrt{{\boldsymbol{\alpha}}_{t}}\mathbf{y}_{t-1},\sqrt{1-{\boldsymbol{\alpha}}_{t}}\mathbf{I})$, which is a Gaussian centered around a copy of $\mathbf{y}_{t-1}$ to which we added noise following a schedule $10<{\boldsymbol{\alpha}}_{1}<{\boldsymbol{\alpha}}_{2}<...<{\boldsymbol{\alpha}}_{T}=$ 1.
@@ 3 Variational Inference With
The goal of DDVI is to fit a latent variable model $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})$. We assume that $p_{\boldsymbol{\theta}}$ factorizes as $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ and admits efficient sampling: examples of such $p_{\boldsymbol{\theta}}$ include Bayesian networks and variational autoencoders (VAEs) (Kingma and Welling 2013).
@@ 3.1 Modeling Family: Diffusion-Based Posteriors
[…]
@@
We define the relationship between $\mathbf{y}$ and $\mathbf{z}$ via a forward diffusion process $r(\mathbf{y}|\mathbf{z},\mathbf{x})=r(\mathbf{y}_{1:T}|\mathbf{z},\mathbf{x})=r(\mathbf{y}_{1}|\mathbf{z},\mathbf{x})\prod_{t=1}^{T-1}r(\mathbf{y}_{t+1}|\mathbf{y}_{t},\mathbf{x})$, which transforms $\mathbf{z}$—the latent whose intractable posterior we seek to approximate—into $\mathbf{y}_{T}$, whose posterior is easier to model (possibly conditioned on $\mathbf{x}$). Examples of $r$ include Gaussian forward diffusion processes and discrete noising processes (Austin et al. 2021). The model $q_{\boldsymbol{\phi}}$ is trained to approximately reverse this forward diffusion process.
@@ 3.2 Learning Objective: A Markovian ELBO
The standard approach to fit auxiliary-variable generative models (Maaløe et al. 2016) is to apply the ELBO twice:
@@ 3.3 Refining the Objective: A Regularized ELBO
Notice that optimizing $\mathcal{L}_{\text{reg}}$ involves sampling from the approximate reverse process $q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$ to match the true reverse process $r(\mathbf{y}|\mathbf{z},\mathbf{x})$: this is the opposite of diffusion training, where we would sample from $r$ to fit $q$. This type of on-policy learning of $q$ has been studied in the context of approximate inference (Zhang and Chen 2021); however, it requires backpropagating through $T$ samples, which may hamper training, and it optimizes a mode-covering divergence that may struggle to fit complex $p(\mathbf{z})$.
@@
We may use the Markov structure in $q,r$ to rewrite $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$ as a sum of $T$ terms, one per Markov step. The derivation is identical to that used to obtain the ELBO of a diffusion model (Sohl-Dickstein et al. 2015), and yields an expression of the same form:
@@
$-\text{KL}(r(\mathbf{y}_{T}|\mathbf{z})||q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x})).$
@@ 3.4 Optimization: Extending Wake-Sleep
We also consider a light-weight algorithm, in which $r(\mathbf{y}|\mathbf{z})$ and $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y})$ do not depend on $\mathbf{x}$. This scenario admits the following optimization procedure:
@@
A common type of noising process compatible with this bound when $\mathbf{z}$ is continuous is Gaussian diffusion, where we define $r(\mathbf{y}_{t}|\mathbf{y}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\sqrt{1-{\boldsymbol{\alpha}}_{t}}\mathbf{y}_{t-1},{\boldsymbol{\alpha}}_{t}\mathbf{I})$ for a suitable schedule $({\boldsymbol{\alpha}}_{t})_{t=1}^{T}$. We then adopt the parameterization $q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})=\mathcal{N}(\mathbf{y}_{t-1};\mu_{\boldsymbol{\phi}}(\mathbf{y}_{t},\mathbf{x},t),\Sigma_{\boldsymbol{\phi}}(\mathbf{y}_{t},\mathbf{x},t))$. It is then common to parameterize $q_{\boldsymbol{\phi}}$ with a noise prediction network ${\boldsymbol{\epsilon}}_{\boldsymbol{\phi}}$ (Ho, Jain, and Abbeel 2020); the sum of KL divergences can be approximated by $\mathbb{E}_{t,{\boldsymbol{\epsilon}}_{t}\sim r(\mathbf{y}_{0},t)}||{\boldsymbol{\epsilon}}_{t}-{\boldsymbol{\epsilon}}_{\boldsymbol{\phi}}(\sqrt{\bar{{\boldsymbol{\alpha}}_{t}}}\mathbf{y}_{0}+\sqrt{1-\bar{{\boldsymbol{\alpha}}_{t}}}{\boldsymbol{\epsilon}}_{t},\mathbf{x},t)||^{2}$. Other extensions include discrete denoising processes (Austin et al. 2021; Sahoo et al. 2024; Schiff et al. 2024). In the wake-sleep setting, we know both endpoints $\mathbf{y}_{T}\sim q(\cdot|\mathbf{x})$ and $\mathbf{y}_{0}=\mathbf{z}$ of the diffusion process, opening the possibility for applying optimal transport techniques (Cuturi 2013; De Bortoli et al. 2021).
@@ 4.1 Semi-Supervised Learning
Following Makhzani et al. (2015), we extend our algorithm to the semi-supervised learning setting where some data points have labels denoted by $l$. We assume the user provides a model of the form $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z},l)=p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z},l)r(\mathbf{y}|\mathbf{z},l)p_{\boldsymbol{\theta}}(\mathbf{z}|l)p(l)$; we set the variational distributions to $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x},\mathbf{y},l),q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x}),q_{\boldsymbol{\phi}}(l|\mathbf{x})$. In this setting, we consider two cases, depending on whether the label is observed (Kingma et al. 2014). We extend Equation (7) to incorporate the label $l$ corresponding to a data point as follows:
@@
When the label $c$ cannot be observed, we treat it as a latent variable and modify the learning objective $\mathcal{U}_{\text{semi}}=\sum_{c}q_{\boldsymbol{\phi}}(l|\mathbf{x})\mathcal{L}_{\text{semi}}(\mathbf{x},l,{\boldsymbol{{\boldsymbol{\theta}}}},{\boldsymbol{\phi}})+\text{KL}(q_{\boldsymbol{\phi}}(l|\mathbf{x})||p(l))$. Therefore, we can conclude a marginal likelihood on our dataset as follows: ~$\tilde{\mathcal{L}}_{\text{semi}}=\sum_{(\mathbf{x},l)\in L}\mathcal{L}_{\text{semi}}(\mathbf{x},l,{\boldsymbol{{\boldsymbol{\theta}}}},{\boldsymbol{{\boldsymbol{\phi}}}})+\sum_{\mathbf{x}\in U}\mathcal{U}_{\text{semi}}(\mathbf{x},{\boldsymbol{{\boldsymbol{\theta}}}},{\boldsymbol{{\boldsymbol{\phi}}}}).$ where $L$ and $U$ are the sets of data with and without labels, respectively.
We also want to guarantee that all model parameters can be learned in all cases, including $q_{\boldsymbol{\phi}}(l|\mathbf{x})$, such that this posterior can be applied as a classifier during inference. Thus, we combine the marginal likelihood with a classification loss to form an extended learning objective: ~$\tilde{\mathcal{L}}_{\text{semi}_{\boldsymbol{\alpha}}}=\tilde{\mathcal{L}}_{\text{semi}}+{\boldsymbol{\alpha}}\cdot\mathbb{E}_{\tilde{p}(\mathbf{x},l)}\left[-\log q_{\boldsymbol{\phi}}(l|\mathbf{x})\right]$

### v5

@@ 1 Introduction
in which we approximate the marginal likelihood $\log\;p_{\boldsymbol{\theta}}(\mathbf{x})$ of a latent variable model $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})$ with an evidence lower bound $\text{ELBO}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ that is a function of an approximate variational posterior $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$. We assume that $p_{\boldsymbol{\theta}}$ factorizes as $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ and admits efficient sampling: examples of such $p_{\boldsymbol{\theta}}$ include Bayesian networks, topic models (blei2003latent), variational autoencoders (VAEs), and broad classes of $p_{\boldsymbol{\theta}}$ defined via modern probabilistic programming frameworks (gordon2014probabilistic).
@@
Maximizing $\text{ELBO}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ over $\boldsymbol{\phi}$} yields a variational posterior $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ that approximates $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$ as well as a tight bound on $\log\;p_{\boldsymbol{\theta}}(\mathbf{x})$ that serves as a learning objective for $p_{\boldsymbol{\theta}}$. The approximation gap $\log\;p_{\boldsymbol{\theta}}(\mathbf{x})-\max_{\boldsymbol{\phi}}\text{ELBO}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ equals precisely $\min_{\boldsymbol{\phi}}\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}))$, which motivates the design of expressive classes of posteriors $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ that reduce this gap. Recent efforts leverage modern generative models—including normalizing flows (rezende2015variational; kingma2016improved) and generative adversarial networks (goodfellow2014generative; makhzani2015adversarial)—as expressive model families for $q_{\boldsymbol{\phi}}$ that tighten the ELBO.
@@
This work seeks to further improve variational inference via expressive posteriors based on diffusion models (ho2020denoising; song2020score). Diffusion methods have become the de-facto standard for high-quality image synthesis (rombach2022high; gokaslan2024commoncanvas). Here, we use diffusion in latent space to parameterize $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$. We train this distribution with a denoising diffusion-like objective that does not involve adversarial training (makhzani2015adversarial) or constrained invertible normalizing flow architectures (kingma2016improved). Samples from $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ are obtained via iterative refinement of $\mathbf{z}$, starting from a Gaussian distribution, and gradually forming one that is multi-modal and complex.
Our work expands upon existing diffusion-based approximate inference methods (berner2022optimal; zhang2021path; vargas2023denoising; zhang2023diffusion; richter2023improved; sendera2024diffusion; akhound2024iterated) that focus on the task of drawing samples from unnormalized distributions $\tilde{p}(\mathbf{z})$ and estimating the partition function $Z=\int_{\mathbf{z}}\tilde{p}(\mathbf{z})d\mathbf{z}$. While these methods are applicable in our setting—we set the unnormalized $\tilde{p}(\mathbf{z})$ to $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})$ such that $Z=p_{\boldsymbol{\theta}}(\mathbf{x})$—they do not make use of characteristics of $p_{\boldsymbol{\theta}}$ that are common in many types of models (VAEs, Bayes networks, etc.), namely the factorization $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ and efficient sampling. We find that leveraging these properties yields simpler algorithms that avoid backpropagating through a sampling process, and that are fast enough to perform learning in addition to inference.
@@
Specifically, we propose denoising diffusion variational inference (DDVI), an approximate inference algorithm defined by a class of approximate posterior distribution based on diffusion and a learning objective inspired by the wake-sleep algorithm (hinton1995wake) that implements regularized variational inference. We also derive extensions of our method to semi-supervised learning and clustering.
@@
We evaluate DDVI on synthetic benchmarks and on a real problem in biological data analysis—inferring human ancestry from genetic data. Our method outperforms strong baselines on 1000 Genomes dataset (siva20081000) and learns a low-dimensional latent space that preserves biologically meaningful structure (haghverdi2015diffusion).
@@ 2 Background
which serves as a tractable surrogate for the marginal log-likelihood (MLL). The gap between the MLL and the ELBO equals precisely $\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}))$—thus, a more expressive $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ may better fit the true posterior and induce a tighter ELBO (kingma2013auto).
Expressive variational posteriors can be formed by choosing more expressive model families—including auxiliary variable methods (maaloe2016auxiliary), MCMC-based methods (salimans2015mcmc), normalizing flows (rezende2015variational)—or improved learning objectives—e.g., adversarial or sample-based losses (makhzani2015adversarial; zhao2017infovae; si2022autoregressive; pmlr-v202-si23a).
The wake-sleep algorithm (hinton1995wake) optimizes an alternative objective
@@
A diffusion model is defined via a user-specified noising process $q$ that maps data $\mathbf{x}_{0}$ into a sequence of $T$ variables $\mathbf{y}_{1:T}=\mathbf{y}_{1},...,\mathbf{y}_{T}$ that represent increasing levels of corruption to $\mathbf{x}_{0}$. We obtain $\mathbf{y}_{1:T}$ by applying a Markov chain $q(\mathbf{y}_{1:T}|\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{y}_{t}|\mathbf{y}_{t-1})$, where we define $\mathbf{y}_{0}=\mathbf{x}_{0}$ for convenience. When $\mathbf{x}_{0}$ is a continuous vector, a standard choice of transition kernel is $q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\sqrt{{\boldsymbol{\alpha}}_{t}}\mathbf{y}_{t-1},\sqrt{1-{\boldsymbol{\alpha}}_{t}}\mathbf{I})$, which is a Gaussian centered around a copy of $\mathbf{y}_{t-1}$ to which we added noise following a schedule $0<{\boldsymbol{\alpha}}_{1}<{\boldsymbol{\alpha}}_{2}<...<{\boldsymbol{\alpha}}_{T}=1$.
@@ 3 Variational Inference With
The goal of DDVI is to fit a latent variable model $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})$. We assume that $p_{\boldsymbol{\theta}}$ factorizes as $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ and admits efficient sampling: examples of such $p_{\boldsymbol{\theta}}$ include Bayesian networks and variational autoencoders (VAEs) (kingma2013auto).
@@ 3.1 Modeling Family: Diffusion-Based Posteriors
DDVI performs variational inference using a family of approximate posteriors $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})=\int_{\mathbf{y}}q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y},\mathbf{x})q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x})d\mathbf{y}$, which themselves contain latent variables $\mathbf{y}\in\mathcal{Y}$. The models $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y},\mathbf{x}),q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x})$ must have tractable densities and support gradient-based optimization over $\boldsymbol{\phi}$}.
@@
We define the relationship between $\mathbf{y}$ and $\mathbf{z}$ via a forward diffusion process $r(\mathbf{y}|\mathbf{z},\mathbf{x})=r(\mathbf{y}_{1:T}|\mathbf{z},\mathbf{x})=r(\mathbf{y}_{1}|\mathbf{z},\mathbf{x})\prod_{t=1}^{T-1}r(\mathbf{y}_{t+1}|\mathbf{y}_{t},\mathbf{x})$, which transforms $\mathbf{z}$—the latent whose intractable posterior we seek to approximate—into $\mathbf{y}_{T}$, whose posterior is easier to model (possibly conditioned on $\mathbf{x}$). Examples of $r$ include Gaussian forward diffusion processes and discrete noising processes (austin2021structured). The model $q_{\boldsymbol{\phi}}$ is trained to approximately reverse this forward diffusion process.
@@ 3.2 Learning Objective: A Markovian ELBO
The standard approach to fit auxiliary-variable generative models (maaloe2016auxiliary) is to apply the ELBO twice:
@@ 3.3 Refining the Objective: A Regularized ELBO
Notice that optimizing $\mathcal{L}_{\text{reg}}$ involves sampling from the approximate reverse process $q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$ to match the true reverse process $r(\mathbf{y}|\mathbf{z},\mathbf{x})$: this is the opposite of diffusion training, where we would sample from $r$ to fit $q$. This type of on-policy learning of $q$ has been studied in the context of approximate inference (zhang2021path); however, it requires backpropagating through $T$ samples, which may hamper training, and it optimizes a mode-covering divergence that may struggle to fit complex $p(\mathbf{z})$.
@@
We may use the Markov structure in $q,r$ to rewrite $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$ as a sum of $T$ terms, one per Markov step. The derivation is identical to that used to obtain the ELBO of a diffusion model (sohl2015deep), and yields an expression of the same form:
@@
$-\text{KL}(r(\mathbf{y}_{T}|\mathbf{z},\mathbf{x})||q_{\boldsymbol{\phi}}(\mathbf{y}_{T}|\mathbf{x}))+\bar{H}(p_{\boldsymbol{\theta}}).$
@@ 3.4 Optimization: Extending Wake-Sleep
We also consider a light-weight algorithm, in which $r(\mathbf{y}|\mathbf{z})$ and $q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z})$ do not depend on $\mathbf{x}$. This scenario admits the following optimization procedure:
@@
[…]
@@ 4.1 Semi-Supervised Learning
Following makhzani2015adversarial, we extend our algorithm to the semi-supervised learning setting where some data points have labels denoted by $l$. We assume the user provides a model of the form $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z},l)=p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z},l)r(\mathbf{y}|\mathbf{z},l)p_{\boldsymbol{\theta}}(\mathbf{z}|l)p(l)$; we set the variational distributions to $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x},\mathbf{y},l),q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x}),q_{\boldsymbol{\phi}}(l|\mathbf{x})$. In this setting, we consider two cases, depending on whether the label is observed (kingma2014semi). We extend Equation (7) to incorporate the label $l$ corresponding to a data point as follows:
@@
When the label $c$ cannot be observed, we treat it as a latent variable and modify the learning objective $\mathcal{U}_{\text{semi}}=\sum_{c}q_{\boldsymbol{\phi}}(l|\mathbf{x})\mathcal{L}_{\text{semi}}(\mathbf{x},l,{\boldsymbol{{\boldsymbol{\theta}}}},{\boldsymbol{\phi}})+\text{KL}(q_{\boldsymbol{\phi}}(l|\mathbf{x})||p(l))$. Therefore, we can conclude a marginal likelihood on our dataset as follows: $\tilde{\mathcal{L}}_{\text{semi}}=\sum_{(\mathbf{x},l)\in L}\mathcal{L}_{\text{semi}}(\mathbf{x},l,{\boldsymbol{{\boldsymbol{\theta}}}},{\boldsymbol{{\boldsymbol{\phi}}}})+\sum_{\mathbf{x}\in U}\mathcal{U}_{\text{semi}}(\mathbf{x},{\boldsymbol{{\boldsymbol{\theta}}}},{\boldsymbol{{\boldsymbol{\phi}}}}).$ where $L$ and $U$ are the sets of data with and without labels, respectively.
We also want to guarantee that all model parameters can be learned in all cases, including $q_{\boldsymbol{\phi}}(l|\mathbf{x})$, such that this posterior can be applied as a classifier during inference. Thus, we combine the marginal likelihood with a classification loss to form an extended learning objective: $\tilde{\mathcal{L}}_{\text{semi}_{\boldsymbol{\alpha}}}=\tilde{\mathcal{L}}_{\text{semi}}+{\boldsymbol{\alpha}}\cdot\mathbb{E}_{\tilde{p}(\mathbf{x},l)}\left[-\log q_{\boldsymbol{\phi}}(l|\mathbf{x})\right]$